OpenStudy (anonymous):

F=kq1q2/r2 solve for k

OpenStudy (anonymous):

Coulombs law :)

OpenStudy (anonymous):

So have you tried solving it yourself yet? :P

OpenStudy (anonymous):

\[F = k \frac{ q_1q_2 }{ r^2 }\] Don't let the variables scare you, it's the same way as you would do it with numbers, what you do to the left side you must also do to the right. So, for example lets say we have the kinematic equation: \[v=v_0+at\] and I ask you to solve for t. What you do on the left side you must also do to the right. Lets start of by getting rid of the initial velocity on the right side (v_0), to do this we have to subtract from both sides so you have: \[v-v_0=\cancel v_0+at-\cancel v_0\] You see it gets cancelled on the right side. Now lets get rid of the acceleration. You can see it's "stuck" on the t which is being multiplied. To remove this, we say alright, what's the opposite operation that will get rid of acceleration on the right side? Division! So we go ahead and do that. \[v-v_0=at \implies \frac{ v-v_0 }{ a } = \frac{\cancel at }{ \cancel a } \implies t = \frac{ v-v_0 }{ a }\]

OpenStudy (anonymous):

Now try solving for k using the same methodology.

OpenStudy (anonymous):

how is that gonna help

OpenStudy (anonymous):

I was showing you an example on how to do algebra, which you should be able to apply to your question, especially if you're using such an equation.

OpenStudy (texaschic101):

f = kq^1q^2 / r^2 --- multiply both sides by r^2 fr^2 = kq^1q^2 -- divide both sides by q^1q^2 (fr^2)/(q^1q^2) = k